### v1

# Optical absorption and luminescence of -LiV2O5 from the Bethe Salpeter Equation

Abstract
-LixV2O5 is obtained by intercalating Li between the layers of V2O5. The partial filling of the split-off conduction band by electron donation from Li leads to significant changes in optical properties. Here we study the electronic band structure of -LiV2O5 using quasiparticle self-consistent (QS) calculations and the optical dielectric function by means of the Bethe Salpeter equation. We find a very strong optical absorption band related to transitions between the filled V- like states to the empty ones with strong polarization along the -direction. We relate this to recent experimental observations of cathodoluminescence (CL) in which a supression of the CL was observed upon addition of Li.
I Introduction
Vanadium pentoxide or V2O5 is a layered oxide with strongly anisotropic properties in the layers thanks to the 1D zigzag chains of edge connected alternatingly up and down pointing square pyramids surrounding each V.[1] Its properties can be significantly tuned by intercalating with Li, Na, Mg and other ions. These atoms are located in between the van der Waals bonded layers in the large open channels between the double zigzag chains, which are connected via a bridge oxygen. See Fig. 1 for the crystal structure. These ions can thus easily diffuse in and out of the structure and this provides a way to tune the properties of the material forming so-called V2O5 bronzes. This makes LixV2O5 attractive as a Li cathode material, since up to a certain concentration, the Li insertion and removal is reversible without significant structural changes. The study of V2O5 as a host for Li cathode battery material goes back to the work of the chemistry Nobel Prize winner Stanley Wittingham [2, 3] and has recently been re-activated by various authors.[4, 5, 6, 7, 8, 9, 10, 11] Optical characterization can play an important role in elucidating the electrochemical insertion and extraction of Li.
Walker et al. [4] used depth resolved Cathodoluminescence (CL) to study the evolution of the CL signal as function of Li uptake and the accompanying changes in the structure of LixV2O5 were also studied by Raman specroscopy by Jarry et al. [5]. The modifications of the optical properties observed in this work can be utilized for electrochromic applications as they modify the color of the crystals. Walker et al. [4] observed a decrease in the lowest energy CL peak (1.8-1.9 eV) upon Li uptake which they analyzed in terms of local density approximation (LDA) band structure calculations by Eyert and Höck [12] and recent X-ray edge spectroscopy studies of LixV2O5 [13, 14]. However, it has recently become clear that the optical properties of V2O5 are not well described by a simple interband transition picture based on band structure theory. Quasiparticle self-consistent calculations by Bhandari et al. [15] showed that the quasiparticle gap is much larger (4.4 eV) than the optical absorption edge of 2.3 eV extracted from Tauc-plots by Kenny and Kannewurf[16]. Recently, this discrepancy was resolved by pointing out the large excitonic binding energies in this material[17, 18]. The interpretation of the optical properties of LixV2O5 thus requires revision in terms of our current understanding of the excitonic properties of V2O5. Polaronic effects are also known to play an important role in V2O5. Since early work on the transport it is known that the electronic transport is essentially of the variable range hopping type associated with small polarons.[19] The strong electron-phonon coupling in V2O5 was shown to lead to self-trapped polarons by Scanlon et al. [20] and later further studied by Ngamwongwan et al. [21] and Watthaisong et al. [22].
In this paper we calculate the electronic structure of LiV2O5 using the QS method[23] and evaluate the optical absorption properties using the recent implementation [24, 25] of the Bethe Salpeter Equation approach based on the linearized-muffin-tin-orbital (LMTO) basis set [26].
First, we note that for one Li per V2O5 unit the lowest split-off conduction band becomes exactly half-filled and this leads to the formation of a magnetic moment and a spin-polarized band strucutre. In analogy with NaV2O5, we expect the spins to be ordered antiferromagnetically along the chains[27]. While the ordering of spins in LiV2O5 requires further study we here adopt this same spin arrangement and study the optical properties in both ferromagnetic (FM) and antiferromagnetic (AFM) LiV2O5. It is known from various studies of the lithiation [10, 11, 28] that upon higher concentration of Li, a phase transition occurs to first the -phase of LiV2O5, which can already accommodate more than one Li per V2O5 unit. For Li2V2O5, the split-off band would become completely filled and magnetic moment formation is then no longer evident. At even higher concentration Li3V2O5 a non-layered (disordered rocksalt-type) -structure has been obtained [11, 10, 28, 29] In the present paper we only consider the regime where the structure stays close to that of -V2O5, i.e. although even there small modifications occur and have been labeled as the - and -phases.
Our primary goal is to provide an alternative interpertation for the CL key observation of a reduction of the lowest CL peak at about 1.8 eV upon the initial lithiation.
II Computational Methods
The band structures are calculated using the QS method [30], for which details of the implementation are given in [23]. This is a variant of Hedinâs method in which is the one-electron Greenâs function and the screened Coulomb interaction [31, 32] and determine the dynamic self-energy . In the quasiparticle self-consistent approach, a non-local but energy independent exchange correlation potential is extracted from the energy-dependent self-energy , with the eigenstates of the non-interacting Hamiltonian and meaning the Hermitian part. This allows to iteratively update and thereby optimize the starting from a density functional theory (DFT) functional such as LDA or generalied gradient approximation (GGA) and makes the final results independent of the starting exchange-correlation functional choice. It focuses on the quasiparticle energies rather than the full self-energy or interacting one-electron Greenâs function and this approach was further justified by Ismail-Beigi[33]. The Questaal code implementation of this method used here uses a muffin-tin-orbital basis set and is described in [26]. For the two-point quantities like the bare and screened Coulomb potentials and the polarization and inverse dielectric function it uses a mixed-product basis set, including plane waves restricted to the interstitial region and products of partial waves times spherical harmonics in the muffin-tin spheres. This idea originates from [34] and provides an efficient way to describe the screening at short range which does not require high energy partial waves or high energy conduction bands. More recently, the Bethe Salpeter Equation (BSE) method [35] which uses the two-particle Greenâs function to study the optical properties incuding local field and electron-hole interaction effects, was also implemented in the Questaal code package and an improvement beyond the random phase approximation (RPA) for the screening of the in also based on the BSE was implemented and described in [24, 25].
Details of the present calculation are similar to our recent study of optical properties in V2O5 [18]. The convergence parameters were carefully optimized. We use a double set of smoothed Hankel function envelope functions with optimized decay lengths and associated smoothing radii with angular momenta for V and O and for Li and include V- semi-core states as local orbitals. We use a k-point mesh and a cut-off of the self-energy of 2.13 Rydberg above which it is approximated by an average diagonal value. We start the QS calculations with a GGA calculation, using the Perdew-Burke-Ernzerhof (PBE) [36] functional with a Hubbard U of 0.1 Rydberg on the V- orbitals to initiate a magnetic moment, but which is then switched off as the self-energy takes over the role of the Hubbard . The results thus do not depend on this initial choice, which is only used to start from a band structure with a spin-splitting due to the half-filling of the narrow spit-off conduction band.
Unlike our previous study on V2O5 which used the QS method in which electron-hole interaction effects are included in the screening of , we here use the original QS method. While this increases the gaps (indirect, lowest direct and direct at ) of V2O5 by about 0.6 eV compared to the more accurate approach, the exciton binding energies are then also larger and the energy of the lowest optical excitations are only overestimated by 0.4 eV. We are here focusing on qualitative changes due to Li and avoided the more time consuming BSE evaluation of screened .
III Results
We start by comparing the QS band structure results for pure V2O5 with those of FM LiV2O5 shown in Fig. 2a-b. As already announced in the introduction, we can see that the narrow split-off conduction band at about 4 eV above the VBM, splits in spin-up and spin down bands, the lower of which is then occupied by the extra electron donated by the Li. The exchange splitting of this band is about 2.9 eV. The splitting between the two lowest majority spin conduction bands is about 2.8 eV at and 2.1 at , where it is lowest. In pure V2O5 at , the split-off band lies about 1.5 eV below the bottom of the main set of conduction bands while this splitting is lower at , where it is about 0.8 eV in V2O5. So, this makes sense: at its maximum near , the split-off band in V2O5 is about 0.8 eV below the main CBM and the spin-splitting induced in LiV2O5 then places the majority spin 1.4 eV below this point, so at about 2.2 eV below the main CBM or majority spin. We may also notice that the highest valence band which in pure V2O5 shows a significant dispersion along becomes flatter in LiV2O5. This indicates that the insertion of Li disrupts the interlayer hopping of the O- like states. The split-off band of majority spin has a band dispersion mostly along the direction, which is the -direction along the chains and has a band width or about 0.8 eV. In FM LiV2O5 at the bottom of the split-off band lies about 2.2 eV above the rather flat VBM of the O- bands. The gap at from the O- like VBM to the majority spin CBM at is 5.26 eV, which is slightly larger than in pure V2O5 where it is 5.07 eV. In a simple band-to-band transition picture,we may now expect new optical transitions between the occupied majority spin band at the lowest empty band of the same spin, at about 2 eV. Next, we can see that in AFM LiV2O5, the situation is essentially the same but now the spin-up and spin down bands are exactly superposed. Note that the spin up bands on one V-O-V rung become the spin-down bands on the next rung along the -direction and vice versa. Because of the doubling of the cell along , the bands are folded. We may note that the higest occupied V- derived bands become almost dispersionless but the lowest gap is similar.
In terms of magnetic properties, we note that the net magnetic moment of the FM cell is 2 . Since there are four V per cell, this means effectively 0.5 per V. Indeed, because there are two formula units per unit cell, there are two bands of majority spin that become exactly filled while their spin-down counterpart stays empty. It indicates that we rather should consider the magnetic moment as being associated with a single V-O-V rung and equally spread over the two V in that rung. This makes sense as the lower split-off band is essentially a molecular type state formed by V- orbitals which is odd versus the mirror plane perpendicular to the -crystallographic axis. As such it cannot interact with the O- states of the bridge oxygen between them, while the higher conduction band at the bottom of the main conduction band manifold is an even combination of the same orbitals which does feel the antibonding -type interaction with the O- of the bridge. The O- and are orthogonal to the orbitals of the two V along the direction and do not come into play. The total energy can at present not reliably be calculated within the QS method. This would require a trace over all states, including very high-lying empty states via the Luttiner-Ward formula. Linear response function calculations indicate that the coupling between spins is antiferromagnetic along the -direction, and this agrees with known experimental data in the related NaV2O5. We reserve a discussion of the magnetic properties to a separate paper. We may then also assume that even in the paramagnetic case, locally magnetic moments form but are disordered and the electronic structure locally is similar to the one we here obtain for the FM case.
Next, we consider the optical properties calculated in BSE and the independent particle approximation (IPA) in Fig. 3. We show the results for both BSE (solid lines) and IPA (dashed lines) for (black, blue) and (red, green). In blue and green are the results using a converged set of 64 valence and 40 conduction bands, i.e. including all O- and V- derived bands, while in black and red we have used only valence and conduction bands. This shows that convergence in the number of bands is important. We do not incude the spectra as this is less relevant being polarized perpendicular to the layers and therefore difficult to measure experimentally. Compared to pure V2O5, [18, 17] where the absorption starts with excitons at about 3.0 eV, the most notable is a new strong absorption band near 2 eV, polarized almost exclusively along . To better see the converged BSE results in the energy region above eV we also show these curves multiplied by a factor 10 in dash-dotted lines and corresponding colors. The new absorption band is already present in the IPA as a narrow peak at 2.45 eV but shifts to 1.85 eV in BSE. At the same time, the charge transfer exciton related to the O- to V- transitions also present in pure V2O5 at about 3 eV also shows up in BSE for both polarizations and the onset of band-to band continuum lies at above 4 eV, corresponding to the qusiparticle gap.
The main new feature peaking at 1.85 eV eV corresponds to the transitions between the occupied split-off band to the lowest unoccupied conduction band, as can be seen in Fig. 4. This feature is similar to what was pointed out in NaV2O5 in Bhandari et al. [27] and was in that material also studied by electron loss spectroscopy[37]. In NaV2O5, experimentally it lies at about 1 eV rather than near 2 eV. However, the QS gap between these states is also near 2 eV similar to the situation in LiV2O5. Thus, strong excitonic effects again influence this optical transition as we can see by comparing BSE with IPA. Closer inspection of the excitonic eigenvalues, shows that there are also nearly dark excitons starting at about 1.4 eV.
A minor difference may result between Na and Li from the difference in distortion of the structure. The angle between the Ob-V- gives a measure of the tilting of the pyramid with its Ov towards the inserted alkali cation and is 102â for Li and 104â for Na. The sharper angle means more tilting. The distance from the vanadyl Ov toward the Na is 2.707 Ă Â while toward the Li is 2.66 Ă . These stronger distortions are consistent with a slightly stronger downward shift of the occupied V- like state by interacting with the intercalated cation.
The reason this absorption is polarized exclusively along is that it corresponds to transitions between odd and even states with respect to the mirror-plane passing through the bridge O and perpendicular to the direction as was pointed out above. Since this is a transition between the molecular localized type states centered on the bridge O and having mainly V- character for both states, is it mainly an intra-atomic transition and thereby significantly stronger than the charge-transfer type transitions. This can be seen in Fig.5. We can clearly see the character of the states. The strongest hole contribution is on the V where the electron is placed and its neighbor across the V-O-V rung.
IV Discussion of the cathodoluminscence
Walker et al[4] studied changes in the electronic structure of LixV2O5 using depth resolved CL. Each of their peaks shows considerable fine structure that evolves with depth from bulk to surface and is also shown to be sensitive to remote oxygen plasma treatment, indicating that O-vacancies play a role. It is here not our intent to interpret these fine changes but rather to understand the broad main features and their changes between pure V2O5 and LiV2O5. Their spectrum for pure V2O5 shows a first peak just below 2 eV with fine structure at 1.8 and 1.9 eV, followed by a smaller peak near 3 eV and merging peaks at 3.6 and 4.1 eV. The main change upon lithiation is that the feature near 2 eV is strongly suppressed and essentially absent but can be partially restored by delithiation. These authors interpreted the feature near 2 eV as resulting from recombination of V- like split-off conduction electrons recombining with holes at the VBM while the higher peaks were associated with other V- orbitals, such as and . In a CL experiment electron hole pairs are first created by a high energy electron impact and then relax to the lowest states near the band edges before they radiatively recombine and emit light. So, first, it is not clear in this interpretation why the observed features are associated with density of states features in the conduction band only while ignoring the valence band structure and assuming that all holes have relaxed to the VBM. This would require one to assume a much faster hole than electron relaxation, for which there is no direct evidence, to the best of our knowledge. Second, the feature below 2 eV is about 0.5 below the aborption onset which was found to be near 2.35 eV by Kenny and Kannewurf[16]. Furthermore, adding Li would partially fill the split-off band and while one could imagine this to suppress absorption to these states, it would not really affect luminescent recombination between the split-off band and the VBM. The main observation of a decrease in the CL peak below 2 eV thus remains unexplained. Here we propose an alternative explanation.
First, luminescence and absorption are proportional to but in luminescence the lower energy region of the spetrum should be emphasized by the carrier relaxation toward the lowest electron and hole states. Furthermore, spontaneous luminescence is proportional to . Second, the optical absorption in pure V2O5 has recently been shown to be dominated by very strongly bound excitons [17, 18]. The first strong absorption occurs at 3 eV followed by peaks at 4 eV and 5-6 eV with the quasiparticle gap at 3.8 eV. We thus associate the CL peaks between 3-4 eV with the direct excitons. The absorption onset of V2O5 near 2.3-2.6 eV on the other hand is still not entirely understood but is likely associated with indirect excitons and/or a dark exciton becoming activated by phonons or disorder as discussed in [38, 18]. But while there can be a Stokes shift between absorption and emission in localized state systems related to different structural relaxation in the ground and excited states, the CL feature at 1.8-1.9 eV is clearly not the same as the onset of absorption. We associate it with a polaronic emission or alternatively a self-trapped exciton. In fact, it has been shown by various calculations [20, 21, 22] that strong electron-phonon coupling leads to the formation of small electron polarons in V2O5. The added electrons are usually associated with O-vacancies, in particular the vanadyl-O-vacancy. Scanlon et al. [20] showed that O-vacancies produce defect states at about 1 eV above the VBM similar to Li interstitials when a DFT calculation is used which favors the formation of localized states and polaronic deformation. This was confirmed by Ngamwongwan et al. [21] and Watthaisong et al. [22] who showed that self-trapping of excitons ccurs near an added electron even without the presence of O-vacancies or nearby Li interstitials. The self-trapped polaron is found near the middle of the gap at about 1 eV above the VBM, while the O-vacancy related states depend on which type of O is removed. We thus propose that the CL feature at 1.8-1.9 eV is associated with recombination of a self-trapped electron with holes at the VBM induced in the CL experiment. We can also view this as a self-trapped indirect exciton. It is not straightforward to extract the energy position of this band from the GGA+U calculations of [20, 21, 22] as all of these miss the strong increase in quasiparticle gap we find in our QS calculations. However, we may assume that the polaron band lies close to the position of our Li induced band in LiV2O5 which is indeed about 2 eV above the VBM. While in our discussion above, this was associated with the spin splitting of the split-off band, the relaxation of the atoms near the Li which tilt the O-pyramids near each V toward the Li play also a decisive role in this feature being shifted down toward the middle of the gap. So, then why is this polaronic emission band reduced upon lithiation? At first one might think that the Li brings in more electrons and hence more polaron like states. However, there is now also a very strong absorption at about the same energy as the emission peak. We therefore propose that the emission is reduced by the strong absorption at this same energy. In other words, we propose that it it is the phenomenon of self-reabsorption that reduces the cathodolumnescence at 1.8 eV.
V Conclusions
In this paper we showed that Li electron donation in -LiV2O5 leads to a splin splitting of the split-off conduction band which creates an occupied band in the middle of the gap. This then leads to strong optical absorption between equal spin bands with intra-atomic V- character. Essentially transitions occur between antisymmetric and symmetric linear combinations of the orbitals on the two V across a V-O-V rung, which form strongly localized molecular type states. This strong absorption is polarized exclusively along the direction. We propose a different interpretation of the CL experiments of Walker et al. [4] with the main peaks at 3-4 eV being associated with the charge transfer direct excitons of V2O5 and the 1.8 eV feature as an electron self-trapped polaron emission which is suppressed in LiV2O5 by self-absorption.