### v5

# Optical absorption and luminescence of -LiV2O5 from the Bethe Salpeter Equation

Abstract
-LixV2O5 is obtained by intercalating Li between the layers of V2O5. The partial filling of the split-off conduction band by electron donation from Li leads to significant changes in optical properties. Here we study the electronic band structure of -LiV2O5 using quasiparticle self-consistent (QS) calculations and the optical dielectric function by means of the Bethe-Salpeter Equation (BSE). The half-filling of the narrow split-off band leads to a spin-splitting and formation of magnetic moments of 0.5 per V which order antiferromagnetically along the chain or -direction. The imaginary part of the dielectric function shows a very strong optical absorption band near 2 eV for polarization along the -direction already in the independent particle approximation but red shifted in the BSE. It is about ten times stronger in intensity than the lowest exciton peaks in pure V2O5 or than the -direction. This absorptions stems from a localized transition between the occupied V- derived band, which is odd with respect to the -mirror plane to the higher lying empty band formed from the same V- orbitals but even with respect to that mirror-plane, which explains its polarization and large oscillator strength. It is found both for the ferromagnetic and antiferromagnetic arrangement of the spins along the -direction. We relate our main finding to a recent experimental study of cathodoluminescence (CL) in which a suppression of the lowest CL peak was observed upon addition of Li. Based on our and literature results, we present a different interpretation of the CL peaks from that study, which was based on the orbital character in a band-to-band picture. The lowest CL peak near 1.8 eV, which lies well below the indirect absorption onset of V2O5 is proposed to be related to recombination of a self-trapped electron polaron, resulting from oxygen vacancies, with a hole at the valence band maximum and is suppressed in LiV2O5 by the strong self–absorption from the Li induced occupied band to the higher empty bands at about the same energy.
I Introduction
Vanadium pentoxide or V2O5 is a layered oxide with strongly anisotropic properties in the layers thanks to the 1D zigzag chains of edge connected alternating up and down pointing square pyramids surrounding each V.[1] Its properties can be significantly tuned by intercalating with Li, Na, Mg and other ions. These atoms are located in between the van der Waals bonded layers in the large open channels between the double zigzag chains, which are connected via a bridge oxygen. See Fig. 1 for the crystal structure. These ions can thus easily diffuse in and out of the structure and this provides a way to tune the properties of the material forming so-called V2O5 bronzes. This makes LixV2O5 attractive as a Li cathode material, since up to a certain concentration, the Li insertion and removal is reversible without significant structural changes. The study of V2O5 as a host for Li cathode battery material goes back to the work of the chemistry Nobel Prize winner Wittingham [2, 3] and has recently been re-activated by various authors.[4, 5, 6, 7, 8, 9, 10, 11] Optical characterization can play an important role in elucidating the electrochemical insertion and extraction of Li.
Walker et al. [4] used depth resolved Cathodoluminescence (CL) to study the evolution of the CL signal as function of Li uptake and the accompanying changes in the structure of LixV2O5 were also studied by Raman spectroscopy by Jarry et al. [5]. The modifications of the optical properties observed in this work can be utilized for electrochromic applications as they modify the color of the crystals. Walker et al. [4] observed a decrease in the lowest energy CL peak (1.8-1.9 eV) upon Li uptake which they analyzed in terms of local density approximation (LDA) band structure calculations by Eyert and Höck [12] and recent X-ray edge spectroscopy studies of LixV2O5 [13, 14]. However, it has recently become clear that the optical properties of V2O5 are not well described by a simple interband transition picture based on band structure theory. Quasiparticle self-consistent calculations by Bhandari et al. [15] showed that the quasiparticle gap is much larger (4.4 eV) than the optical absorption edge of 2.3 eV extracted from Tauc plots by Kenny and Kannewurf[16]. Recently, this discrepancy was resolved by pointing out the large excitonic binding energies in this material[17, 18]. The interpretation of the optical properties of LixV2O5 thus requires revision in terms of our current understanding of the excitonic properties of V2O5. Polaronic effects are also known to play an important role in V2O5. Since early work on the transport it is known that the electronic transport is essentially of the variable range hopping type associated with small polarons.[19] The strong electron-phonon coupling in V2O5 was shown to lead to self-trapped polarons by Scanlon et al. [20] and later further studied by Ngamwongwan et al. [21] and Watthaisong et al. [22].
In this paper we calculate the electronic structure of LiV2O5 using the QS method[23] and evaluate the optical absorption properties using the recent implementation [24, 25] of the Bethe-Salpeter Equation approach based on the linearized-muffin-tin-orbital (LMTO) basis set [26].
First, we note that for one Li per V2O5 unit the lowest split-off conduction band becomes exactly half-filled and this leads to the formation of a magnetic moment and a spin-polarized band structure. In analogy with NaV2O5, we find the spins to be ordered antiferromagnetically along the chains[27]. While the magnetic properties in LiV2O5 deserve a further study we here adopt this same spin arrangement and study the optical properties in both ferromagnetic (FM) and antiferromagnetic (AFM) LiV2O5. It is known from various studies of the lithiation [10, 11, 28] that upon higher concentration of Li, a phase transition occurs to first the -phase of LiV2O5, which can already accommodate more than one Li per V2O5 unit. For Li2V2O5, the split-off band would become completely filled and magnetic moment formation is then no longer evident. At even higher concentration Li3V2O5 a non-layered (disordered rocksalt-type) -structure has been obtained [11, 10, 28, 29] In the present paper we only consider the regime where the structure stays close to that of -V2O5, i.e. although even there small modifications occur and have been labeled as the - and -phases.
Our primary goal is to provide an alternative interpretation for the CL key observation of a reduction of the lowest CL peak at about 1.8 eV upon the initial lithiation.
II Computational Methods
The band structures are calculated using the QS method [30], for which details of the implementation are given in [23]. This is a variant of Hedin’s method in which is the one-electron Green’s function and the screened Coulomb interaction [31, 32] and determine the dynamic self-energy . In the quasiparticle self-consistent approach, a non-local but energy independent exchange correlation potential is extracted from the energy-dependent self-energy , with the eigenstates of the non-interacting Hamiltonian and meaning the Hermitian part. This allows to iteratively update and thereby optimize the starting from a density functional theory (DFT) functional such as LDA or generalized gradient approximation (GGA) and makes the final results independent of the starting exchange-correlation functional choice. It focuses on the quasiparticle energies rather than the full self-energy or interacting one-electron Green’s function and this approach was further justified by Ismail-Beigi[33].
The Questaal code implementation of this method used here uses a muffin-tin-orbital basis set and is described in [26]. For the two-point quantities like the bare and screened Coulomb potentials and the polarization and inverse dielectric function it uses a mixed-product basis set, including plane waves restricted to the interstitial region and products of partial waves times spherical harmonics in the muffin-tin spheres. This idea originates from [34] and provides an efficient way to describe the screening at short range which does not require high energy partial waves or high energy conduction bands. More recently, the Bethe Salpeter Equation (BSE) method [35] which uses the two-particle Green’s function to study the optical properties including local field and electron-hole interaction effects, was also implemented in the Questaal code package as described in [24, 25]. In these papers, the BSE method is also used to include electron-hole interactions in the screening of by including ladder-diagrams and which is then called QS. This reduces the QS gap slightly but also decreases the exciton binding energies correspondingly. Unlike our previous study on V2O5 which used the QS method, we here use the original QS method with RPA screening for because the QS method becomes rather expensive for the large 64 atom cells needed for the antiferromagnetic structure. While this increases the gaps (indirect, lowest direct and direct at ) of V2O5 by about 0.6 eV compared to the more accurate approach, the exciton binding energies are then also larger and the energy of the lowest optical excitations are only overestimated by 0.4 eV. We are here focusing on qualitative changes due to Li and avoided the more time consuming BSE evaluation of screened .
Details of the present calculation are similar to our recent study of optical properties in V2O5 [18]. The convergence parameters were carefully optimized. We use a double set of smoothed Hankel function envelope functions with optimized decay lengths and associated smoothing radii with angular momenta for V and O and for Li and include V- semi-core states as local orbitals. We use a well converged k-point mesh for the QS and a cut-off of the self-energy of 2.13 Rydberg above which it is approximated by an average diagonal value. The product basis set uses a default angular momentum cut-off of and a maximum reciprocal lattice vector for the interstitial plane wave part (the GCUTX parameter) of 2.6 Ry1/2, while for the muffin-tin orbitals in the interstitial region in the calculation, a cut-off (named GCUTB) of 3.2 Ry1/2 is used. Because of the about 3 times larger unit cell lattice constant than the and lattice constants in the V2O5 cell and the small dispersion of the bands in the direction perpendicular to the layers, we use a smaller number of divisions in the reciprocal direction. Comparing a and a mesh we find that the gap in FM LiV2O5 case differs by less than 0.1 eV. The results shown in the figure correspond to the mesh which is adequately converged. Note that a finer mesh such as is used for the DFT self-consistency and the self-energy is interpolated to this mesh in the course of the QS iterations. A similar procedure is used to obtain the bands along the symmetry lines by Fourier transforming to a real space representation and back to the k-points along the symmetry lines. For the AFM cell which is doubled in the direction, our final calculations were done with a mesh. We start the QS calculations with a GGA calculation, using the Perdew-Burke-Ernzerhof (PBE) [36] functional with a Hubbard of 0.1 Rydberg on the V- orbitals to initiate a magnetic moment, but which is then switched off as the self-energy takes over the role of the Hubbard . The results thus do not depend on this initial choice, which is only used to start from a band structure with a spin-splitting due to the half-filling of the narrow split-off conduction band. For the BSE calculations we use a similar k-mesh. While BSE for Wannier type excitons typically requires a fine mesh near the band extrema, this is not the case here because the band edges are very flat and the excitons are quite localized Frenkel type excitons.
III Results
We start by comparing the QS band structure results for pure V2O5 with those of FM LiV2O5 shown in Fig. 2a-b. The atomic orbital character of the bands in LiV2O5 is similar to that in V2O5 as discussed in [15] with the Li levels high in the conduction band and merely donating their electron to the lowest V- bands. The manifold of occupied bands shown here consists of the O- derived bands, and the conduction bands are mostly V- derived with higher antibonding O- character as we go higher in energy. As already announced in the introduction, we can see that the narrow split-off conduction band at about 4 eV above the VBM, splits in spin-up and spin down bands, the lower of which is then occupied by the extra electron donated by the Li. The exchange splitting of this band is about 2.9 eV. This is the splitting between the red and green bands in Fig.2b. The splitting between the lowest empty and the highest occupied majority spin bands is about 2.8 eV at and 2.1 at , where it is lowest. This splitting is relevant to the optical transitions between equal spin bands. In the calculation, this gap is about 1.5 eV so, the self-consistency of QS is important. In pure V2O5 at , the split-off band lies about 1.5 eV below the bottom of the main set of conduction bands while this splitting is lower at , where it is about 0.8 eV in V2O5. So, this makes sense: at its maximum near , the split-off band in V2O5 is about 0.8 eV below the main CBM and the spin-splitting induced in LiV2O5 then places the majority spin at half the spin splitting (1.4 eV) below this point, so at about 2.2 eV below the main CBM of majority spin.
We may also notice that the highest valence band which in pure V2O5 shows a significant dispersion along becomes flatter in LiV2O5. This indicates that the insertion of Li disrupts the interlayer hopping of the O- like states. The split-off band of majority spin has a band dispersion mostly along the direction, which is the -direction along the chains and has a band width of about 0.8 eV. In FM LiV2O5 at the bottom of the split-off band lies about 2.2 eV above the rather flat VBM of the O- bands. The gap at from the O- like VBM to the majority spin CBM at is 5.26 eV, which is slightly larger than in pure V2O5 where it is 5.07 eV. In a simple band-to-band transition picture, we may now expect new optical transitions from the occupied majority spin band to the lowest empty band of the same spin, at about 2 eV.
Next, we can see that in AFM LiV2O5, the situation is essentially the same but now the spin-up and spin down bands are exactly superposed. Note that the spin up bands on one V-O-V rung become the spin-down bands on the next rung along the -direction and vice versa. Because of the doubling of the cell along , the bands are folded in two along . We may note that the highest occupied V- derived bands become almost dispersion-less but the lowest gap is similar. The band gap is 2.337 eV in QS, 1.846 eV in and 0.728 eV in the initial GGA+U calculation.
In terms of magnetic properties, we note that the net magnetic moment of the FM cell is 2 . Since there are four V per cell, this means effectively 0.5 per V. Indeed, because there are two formula units per unit cell, there are two bands of majority spin that become exactly filled while their spin-down counterpart stays empty. It indicates that we rather should consider the magnetic moment as being associated with a single V-O-V rung and equally spread over the two V in that rung. This makes sense as the lower split-off band is essentially a molecular type state formed by V- orbitals which is odd versus the mirror plane perpendicular to the -crystallographic axis. As such, it cannot interact with the O- states of the bridge oxygen between them, while the higher conduction band at the bottom of the main conduction band manifold is an even combination of the same orbitals which does feel the antibonding -type interaction with the O- of the bridge. The O- and are orthogonal to the orbitals of the two V along the direction and do not come into play. The AFM state is found to have lower energy than the FM state and therefore we focus in the remainder on the AFM configuration. Linear response function calculations indicate that the coupling between spins is antiferromagnetic along the -direction, and this agrees with known experimental data in the related NaV2O5 and our findings here for LiV2O5. We reserve a full discussion of the magnetic properties, such as the exchange interactions, spin waves and Néel temperature for a separate paper. We assume that even in the paramagnetic case, local magnetic moments form but are disordered and the electronic structure locally is similar to the one we here obtain for the FM case.
Next, we consider the optical properties calculated in BSE and the independent particle approximation (IPA) in Fig. 3 compared to those of pure V2O5. We note that for there is a very strong absorption band near 2 eV, which is already prominent in the IPA but as usual red-shifted and modified in shape with two peaks in the BSE. To allow for comparison with pure V2O5 we show it on a different vertical scale in the middle panel. This allows us to better see the energy range above 3 eV. We may note that apart from a upward shift of about 0.6 eV the spectrum in this range is similar to that of pure V2O5. This shift is related to the flattening of the VBM in LiV2O5 noted earlier for the band structures. For , the spectrum shows a much smaller absorption near 2 eV and at higher energies the spectrum resembles that of pure V2O5.
In Fig 4 we show the optical dielectric function for the FM case. We see a similar strong absorption band near 2 eV for the polarization. These calculations include 32 valence bands (VB) and 20 conduction bands (CB) for the FM cell, 64 VB and 40 CB for the AFM cell, which contains twice the number of atoms, and 30 VB, 20 CB for V2O5. Including all the O- derived and V- bands is important for the convergence of these spectra.
The main new feature, which has its first peak more precisely at 1.85 eV, corresponds to the transitions between the occupied split-off band to the lowest unoccupied conduction band, as can be seen in Fig. 5. This feature is similar to what was pointed out in NaV2O5 in Bhandari et al. [27] and was in that material also studied by electron loss spectroscopy[37]. In NaV2O5, experimentally it lies at about 1 eV rather than near 2 eV. However, the QS gap between these states is also near 2 eV similar to the situation in LiV2O5. Thus, strong excitonic effects again influence this optical transition as we can see by comparing BSE with IPA. Closer inspection of the excitonic eigenvalues, shows that there are also nearly dark excitons starting at about 1.4 eV.
A minor difference may result between Na and Li from the difference in distortion of the structure. The angle between the Ob-V- gives a measure of the tilting of the pyramid with its Ov towards the inserted alkali cation and is 102∘ for Li and 104∘ for Na. The sharper angle means more tilting. The distance from the vanadyl Ov toward the Na is 2.707 Å while toward the Li is 2.66 Å. These stronger distortions are consistent with a slightly stronger downward shift of the occupied V- like state by interacting with the intercalated cation.
The reason this absorption is polarized almost exclusively along is that it corresponds to transitions between odd and even states with respect to the mirror-plane passing through the bridge O and perpendicular to the direction as was pointed out above and explained in more detail in [15, 27] Since this is a transition between the molecular localized type states centered on the bridge O and having mainly V- character for both states, but with a O- antibonding character for the upper empty band, it has a strong intra-atomic (on V) as well as localized character confined to the rung of the ladder and is therefore significantly stronger than the charge-transfer type transitions between the main O- VBM and V- like CBM in pure V2O5. This can be seen in Fig.6. We can clearly see the character of the states. The strongest hole contribution is on the V where the electron is placed and its neighbor across the V-O-V rung.
The very strong peak in in the -direction, also leads to a strong enhancement of the in the BSE and in the IPA in the AFM. These are the static values but including only electronic screening without contributions from the phonons, commonly referred to as . This compares with a value of 6.2 in the -direction in BSE and in the IPA. In pure V2O5, the electronic screening only dielectric constant were calculated using a different approach in [38] and are 6.54 in and 6.08 in the directions. This is a remarkable anisotropy in the dielectric constant.
IV Discussion of the cathodoluminscence
While our main results pertain to the absorption, we here wish to comment on the related cathodoluminescene (CL) spectra, which are the main experimental data on optical properties available at this time. Walker et al[4] studied changes in the electronic structure of LixV2O5 using depth resolved CL. Each of their peaks shows considerable fine structure that evolves with depth from bulk to surface and is also shown to be sensitive to remote oxygen plasma treatment, indicating that O-vacancies play a role. It is here not our intent to interpret these fine changes but rather to understand the broad main features and their changes between pure V2O5 and LiV2O5. Their spectrum for pure V2O5 , which is here reproduced approximately in Fig. 7 by digitizing the data from their figure, shows a first peak just below 2 eV with fine structure at 1.8 and 1.9 eV, followed by a smaller peak near 3 eV and merging peaks at 3.6 and 4.1 eV. The main change upon lithiation is that the feature near 2 eV is strongly suppressed and essentially absent but can be partially restored by delithiation. These authors interpreted the feature near 2 eV as resulting from recombination of V- like split-off conduction electrons recombining with holes at the VBM while the higher peaks were associated with other V- orbitals, such as and .
Considering first Walker et al. ’s interpretation, it is not clear why the observed features are associated with density of states features in the conduction band only while ignoring the valence band structure and assuming that all holes have relaxed to the VBM. This would require one to assume a much faster hole than electron relaxation, for which there is no direct evidence, to the best of our knowledge. Second, the feature below 2 eV is about 0.5 below the indirect absorption onset which was found to be near 2.35 eV by Kenny and Kannewurf[16]. It is thus clearly not a pure band to band transition or related band gap exciton but may be related to defects. Furthermore, adding Li would partially fill the split-off band and while one could imagine this to suppress absorption to these states, it would not really affect luminescent recombination between the split-off band and the VBM. The main observation of a decrease in the CL peak below 2 eV thus remains unexplained. Here we propose an alternative explanation.
Unfortunately, the interpretation of CL spectra is rather complex and a fully first-principles approach is not yet available. In a CL experiment electron hole pairs are first created by a high energy electron impact and then the electrons and holes relax to their lowest states near the band edges by multiple phonon emissions before they radiatively recombine and emit light. That last process is also in competition with non-radiative recombination. The spontaneous luminescence rate arising from the perfect crystal energy bands bands is proportional to the corresponding absorption coefficient with and . Here, is the real part of the index of refraction and is Boltzmann’s constant. However, as the excited electrons relax they increasingly populate lower energy states and hence for a defective crystal CL will more prominently show recombination from the defect levels in the gap to which the excited electrons trickle down.
The optical absorption in pure V2O5 has recently been shown to be dominated by very strongly bound excitons [17, 18]. The first strong absorption occurs at 3 eV followed by peaks at 4 eV and 5-6 eV with the quasiparticle gap at eV in the QS method used in [18] and at slightly higher energies eV for the quasiparticle gap and 3.4 eV for the lowest exciton in QS as seen in Fig. 3. We thus associate the CL peaks between 3-4 eV with the direct excitons and band gap emission. The absorption onset of V2O5 near 2.3-2.6 eV on the other hand is still not entirely understood but is likely associated with indirect excitons and/or a dark exciton becoming activated by phonons or disorder as discussed in [39, 18]. The CL feature at 1.8-1.9 eV is clearly well below the onset of absorption. Thus, it must rather be related to recombination from a defect level in the gap. The most prevalent defect levels in V2O5 are related to oxygen vacancies and self-trapped polarons. In fact both correspond to electrons trapped on a single V. We therefore tentatively associate the lowest CL peak with a emission from a self-trapped polaron recombining with a hole in the VB, which could alternatively be described as a self-trapped exciton. In fact, it has been shown by various calculations [20, 21, 22] that strong electron-phonon coupling leads to the formation of small electron polarons in V2O5. The added electrons are usually associated with O-vacancies, in particular the vanadyl-O-vacancy. Scanlon et al. [20] showed that O-vacancies produce defect states at about 1 eV above the VBM similar to Li interstitials when a DFT calculation is used which favors the formation of localized states and polaronic deformation. This was confirmed by Ngamwongwan et al. [21] and Watthaisong et al. [22] who showed that self-trapping of excitons occurs near an added electron even without the presence of O-vacancies or nearby Li interstitials. In their calculations, the self-trapped polaron is found near the middle of the gap at about 1 eV above the VBM, while the O-vacancy related states depend on which type of O is removed. It is not straightforward to extract the energy position of this band from the GGA+U calculations of [20, 21, 22] as all of these miss the strong increase in quasiparticle gap we find in our QS calculations. However, we may assume that the polaron band lies close to the position of our Li induced occupied band in LiV2O5 since this also corresponds to adding an electron which localizes on a V- state, although it would in the present case be spread over two V. It is indeed about 2 eV above the VBM. Localization on a single V would pull it slightly further down. While in our discussion above, this was associated with the spin splitting of the split-off band, the relaxation of the atoms near the Li which tilt the O-pyramids near each V toward the Li play also a decisive role in this feature being shifted down toward the middle of the gap. In summary, we associate the eV CL band with a transition from this mid gap level to the VBM.
So, then why is this polaronic emission band reduced upon lithiation? At first one might think that the Li brings in more electrons and hence more polaron-like states. However, as we showed in the main results of the previous section there is now also a very strong absorption at about the same energy as the polaron emission peak. While in V2O5 even with some defects, the absorption from the defect levels to the conduction band is negligible compared to the emission resulting from recombination with the VBM holes, this is no longer the case in LiV2O5, where we found as main result a very strong optical absorption peak also at 1.8 eV. In fact, it is about a factor 10 stronger than the lowest exciton peaks in V2O5 and this is due to the more localized character of this optical transition between occupied and empty bands both dominated by the V- orbitals compared to the transition from a V- state to a more delocalized O- VB state involved in the emission process. We therefore propose that the emission is reduced by the strong absorption at this same energy. In other words, we propose that it is the phenomenon of “self-reabsorption” that reduces the cathodoluminescence at 1.8 eV. Interestingly, the self-trapped polaron level lies at about the same energy as the Li induced occupied band and this level lies almost exactly mid-gap so that absorption from this level to the CBM occurs at almost the same energy as recombination with a hole from the VBM.
V Conclusions
In this paper we showed that Li electron donation in -LiV2O5 leads to a spin splitting of the split-off conduction band which creates an occupied band in the middle of the gap. This then leads to strong optical absorption between equal spin bands with intra-atomic V- character. Essentially transitions occur between antisymmetric and symmetric linear combinations of the orbitals on the two V across a V-O-V rung, which form strongly localized molecular type states. This strong absorption is polarized exclusively along the direction. We propose a different interpretation of the CL experiments of Walker et al. [4] with the main peaks at 3-4 eV being associated with the charge transfer direct excitons of V2O5 and the 1.8 eV feature as an electron self-trapped polaron emission which is suppressed in LiV2O5 by self-absorption.
Acknowledgements.
This work was supported by the U.S. Department of Energy Basic Energy Sciences (DOE-BES) under Grant No. DE-SC0008933. Calculations made use of the High-Performance Computing Resource in the Core Facility for Advanced Research Computing at Case Western Reserve University and the Ohio Supercomputer Center. Data Availability The data that support the findings of this article are openly available[40].References
- Sucharitakul et al. [2017] S. Sucharitakul, G. Ye, W. R. L. Lambrecht, C. Bhandari, A. Gross, R. He, H. Poelman, and X. P. A. Gao, V2O5: A 2D van der Waals Oxide with Strong In-Plane Electrical and Optical Anisotropy, ACS Applied Materials & Interfaces 9, 23949 (2017), pMID: 28677951.
- Whittingham [1976] M. S. Whittingham, The role of ternary phases in cathode reactions, Journal of The Electrochemical Society 123, 315 (1976).
- Zavalij and Whittingham [1999] P. Y. Zavalij and M. S. Whittingham, Structural chemistry of vanadium oxides with open frameworks, Acta Crystallographica Section B 55, 627 (1999).
- Walker et al. [2020] M. J. Walker, A. Jarry, N. Pronin, J. Ballard, G. W. Rubloff, and L. J. Brillson, Nanoscale depth and lithiation dependence of v2o5 band structure by cathodoluminescence spectroscopy, J. Mater. Chem. A 8, 11800 (2020).
- Jarry et al. [2020] A. Jarry, M. Walker, S. Theodoru, L. J. Brillson, and G. W. Rubloff, Elucidating Structural Transformations in LixV2O5 Electrochromic Thin Films by Multimodal Spectroscopies, Chemistry of Materials 32, 7226 (2020).
- Levy et al. [2023] Z. Levy, V. C. Ferrari, P. Rosas, M. J. Walker, K. Duddella, M. Haseman, D. Stewart, G. Rubloff, and L. J. Brillson, Lithium Spatial Distribution and Split-Off Electronic Bands at Nanoscale V2O5/LiPON Interfaces, ACS Applied Energy Materials 6, 4538 (2023).
- Warecki et al. [2024] Z. Warecki, V. C. Ferrari, D. A. Robinson, J. D. Sugar, J. Lee, A. V. Ievlev, N. S. Kim, D. M. Stewart, S. B. Lee, P. Albertus, G. Rubloff, and A. A. Talin, Simultaneous Solid Electrolyte Deposition and Cathode Lithiation for Thin Film Batteries and Lithium Iontronic Devices, ACS Energy Letters 9, 2065 (2024).
- Gillette et al. [2016] E. I. Gillette, N. Kim, G. W. Rubloff, and S. B. Lee, Interconnected mesoporous V2O5 electrode: impact on lithium ion insertion rate, Phys. Chem. Chem. Phys. 18, 30605 (2016).
- Chen et al. [2013] X. Chen, E. Pomerantseva, K. Gregorczyk, R. Ghodssi, and G. Rubloff, Cathodic ALD V2O5 thin films for high-rate electrochemical energy storage, RSC Adv. 3, 4294 (2013).
- Yue and Liang [2017] Y. Yue and H. Liang, Micro- and Nano-Structured Vanadium Pentoxide (V2O5) for Electrodes of Lithium-Ion Batteries, Advanced Energy Materials 7, 1602545 (2017).
- Rocquefelte et al. [2003] X. Rocquefelte, F. Boucher, P. Gressier, and G. Ouvrard, First-principle study of the intercalation process in the lixv2o5 system, Chemistry of Materials 15, 1812 (2003).
- Eyert and Höck [1998] V. Eyert and K.-H. Höck, Electronic structure of : Role of octahedral deformations, Phys. Rev. B 57, 12727 (1998).
- De Jesus et al. [2016] L. R. De Jesus, G. A. Horrocks, Y. Liang, A. Parija, C. Jaye, L. Wangoh, J. Wang, D. A. Fischer, L. F. J. Piper, D. Prendergast, and S. Banerjee, Mapping polaronic states and lithiation gradients in individual V2O5 nanowires, Nature Communications 7, 12022 (2016).
- Horrocks et al. [2016] G. A. Horrocks, E. J. Braham, Y. Liang, L. R. De Jesus, J. Jude, J. M. Velázquez, D. Prendergast, and S. Banerjee, Vanadium K-Edge X-ray Absorption Spectroscopy as a Probe of the Heterogeneous Lithiation of V2O5: First-Principles Modeling and Principal Component Analysis, The Journal of Physical Chemistry C 120, 23922 (2016).
- Bhandari et al. [2015] C. Bhandari, W. R. L. Lambrecht, and M. van Schilfgaarde, Quasiparticle self-consistent calculations of the electronic band structure of bulk and monolayer , Phys. Rev. B 91, 125116 (2015).
- Kenny et al. [1966] N. Kenny, C. Kannewurf, and D. Whitmore, Optical absorption coefficients of vanadium pentoxide single crystals, Journal of Physics and Chemistry of Solids 27, 1237 (1966).
- Gorelov et al. [2022] V. Gorelov, L. Reining, M. Feneberg, R. Goldhahn, A. Schleife, W. R. L. Lambrecht, and M. Gatti, Delocalization of dark and bright excitons in flat-band materials and the optical properties of V2O5, npj Computational Materials 8, 94 (2022).
- Garcia et al. [2024] C. Garcia, S. K. Radha, S. Acharya, and W. R. L. Lambrecht, Quasiparticle band structure and excitonic optical response in bulk and monolayer, Phys. Rev. B 110, 085102 (2024).
- Ioffe and Patrina [1970] V. A. Ioffe and I. B. Patrina, Comparison of the Small-Polaron Theory with the Experimental Data of Current Transport in V2O5, Physica Status Solidi (b) 40, 389 (1970).
- Scanlon et al. [2008] D. O. Scanlon, A. Walsh, B. J. Morgan, and G. W. Watson, An ab initio Study of Reduction of V2O5 through the Formation of Oxygen Vacancies and Li Intercalation, The Journal of Physical Chemistry C 112, 9903 (2008).
- Ngamwongwan et al. [2021] L. Ngamwongwan, I. Fongkaew, S. Jungthawan, P. Hirunsit, S. Limpijumnong, and S. Suthirakun, Electronic and thermodynamic properties of native point defects in V2O5: a first-principles study, Phys. Chem. Chem. Phys. 23, 11374 (2021).
- Watthaisong et al. [2019] P. Watthaisong, S. Jungthawan, P. Hirunsit, and S. Suthirakun, Transport properties of electron small polarons in a V2O5 cathode of Li-ion batteries: a computational study, RSC Adv. 9, 19483 (2019).
- Kotani et al. [2007] T. Kotani, M. van Schilfgaarde, and S. V. Faleev, Quasiparticle self-consistent GW method: A basis for the independent-particle approximation, Phys.Rev. B 76, 165106 (2007).
- Cunningham et al. [2018] B. Cunningham, M. Grüning, P. Azarhoosh, D. Pashov, and M. van Schilfgaarde, Effect of ladder diagrams on optical absorption spectra in a quasiparticle self-consistent framework, Phys. Rev. Materials 2, 034603 (2018).
- Cunningham et al. [2023] B. Cunningham, M. Grüning, D. Pashov, and M. van Schilfgaarde, : Quasiparticle self-consistent with ladder diagrams in , Phys. Rev. B 108, 165104 (2023).
- Pashov et al. [2019] D. Pashov, S. Acharya, W. R. Lambrecht, J. Jackson, K. D. Belashchenko, A. Chantis, F. Jamet, and M. van Schilfgaarde, Questaal: A package of electronic structure methods based on the linear muffin-tin orbital technique, Computer Physics Communications , 107065 (2019).
- Bhandari and Lambrecht [2015] C. Bhandari and W. R. L. Lambrecht, Electronic and magnetic properties of electron-doped and , Phys. Rev. B 92, 125133 (2015).
- Delmas et al. [1994] C. Delmas, H. Cognac-Auradou, J. Cocciantelli, M. Ménétrier, and J. Doumerc, The LixV2O5 system: An overview of the structure modifications induced by the lithium intercalation, Solid State Ionics 69, 257 (1994).
- Galy [1992] J. Galy, Vanadium pentoxide and vanadium oxide bronzes—Structural chemistry of single (S) and double (D) layer MxV2O5 phases, Journal of Solid State Chemistry 100, 229 (1992).
- van Schilfgaarde et al. [2006] M. van Schilfgaarde, T. Kotani, and S. Faleev, Quasiparticle Self-Consistent Theory, Phys. Rev. Lett. 96, 226402 (2006).
- Hedin [1965] L. Hedin, New method for calculating the one-particle green’s function with application to the electron-gas problem, Phys. Rev. 139, A796 (1965).
- Hedin and Lundqvist [1969] L. Hedin and S. Lundqvist, Effects of electron-electron and electron-phonon interactions on the one-electron states of solids, in Solid State Physics, Advanced in Research and Applications, Vol. 23, edited by F. Seitz, D. Turnbull, and H. Ehrenreich (Academic Press, New York, 1969) pp. 1–181.
- Ismail-Beigi [2017] S. Ismail-Beigi, Justifying quasiparticle self-consistent schemes via gradient optimization in baym–kadanoff theory, Journal of Physics: Condensed Matter 29, 385501 (2017).
- Aryasetiawan and Gunnarsson [1994] F. Aryasetiawan and O. Gunnarsson, Product-basis method for calculating dielectric matrices, Phys. Rev. B 49, 16214 (1994).
- Onida et al. [2002] G. Onida, L. Reining, and A. Rubio, Electronic excitations: density-functional versus many-body Green’s-function approaches, Rev. Mod. Phys. 74, 601 (2002).
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized Gradient Approximation Made Simple, Phys. Rev. Lett. 77, 3865 (1996).
- Atzkern et al. [2001] S. Atzkern, M. Knupfer, M. S. Golden, J. Fink, A. N. Yaresko, V. N. Antonov, A. Hübsch, C. Waidacher, K. W. Becker, W. von der Linden, G. Obermeier, and S. Horn, One-dimensional dynamics of the d electrons in , Phys. Rev. B 63, 165113 (2001).
- Bhandari and Lambrecht [2014] C. Bhandari and W. R. L. Lambrecht, Phonons and related spectra in bulk and monolayer , Phys. Rev. B 89, 045109 (2014).
- Gorelov et al. [2023] V. Gorelov, L. Reining, W. R. L. Lambrecht, and M. Gatti, Robustness of electronic screening effects in electron spectroscopies: Example of , Phys. Rev. B 107, 075101 (2023).
- [40] https://github.com/Electronic-Structure-Group/LiV2O5_Optics.